OpenStudy (anonymous):

Need help on a Algebra 2 Honors Question!

OpenStudy (anonymous):

4.f(x)= -1 if x ≤ -6 5 if -6 < x ≤ -2 2x + 9 if x > -2 (1 Point)

OpenStudy (anonymous):

I also have part of a table filled out

OpenStudy (anonymous):

I just don't understand what I am supposed to do for the first 4 columns.

OpenStudy (anonymous):

if i was good at this ill help but me im not that good

OpenStudy (anonymous):

I just know about the las column never did any like the first two. I could post other I have already done.

OpenStudy (anonymous):

would that help you help me

OpenStudy (anonymous):

like what r u suppose to do?

OpenStudy (anonymous):

plug in numbers for coordinates and graph it.

OpenStudy (anonymous):

Example fx=2x+9 if x > -2

OpenStudy (anonymous):

Next you would chose a value for x that is greater than -2 which is 0

OpenStudy (anonymous):

oh i have no idea but i can ask my bro to help u

OpenStudy (anonymous):

so now its 2(0)+9 which gives us 9 if we solve it and that is our y value

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

giving us the coordinates for a point at (0,9) if we do a few more we get a graph for two of the columns

OpenStudy (anonymous):

I just don't now what to do if there is no x value for the equation.

OpenStudy (anonymous):

yeahh im confused on the same thin ur talkin bout o dont get it either

OpenStudy (blacksteel):

This function works just like any other in the other two sections. For -6 < x < -2, the function is just y = 5, so for any x value in that range you'll get 5 as your result. Similarly, for any x < -6, y = 1.

OpenStudy (anonymous):

so any number would be -1 for the first two columns

OpenStudy (blacksteel):

For the first two columns, you'll want to use different x values, but the y value will always be -1. So for example, the first two columns might look like: \[\left[ -6 | -1 \right]\]\[\left[ -7 | -1 \right]\]\[\left[ -8 | -1 \right]\]

OpenStudy (anonymous):

and the graph would be a horizontal line on y=-1 right

OpenStudy (blacksteel):

Yes.

OpenStudy (anonymous):

thanks a lot I relly needed help understanding this and this is the only question that did those with nothing in the lesson that talked about it. Thanks again

OpenStudy (blacksteel):

No problem, feel free to message me if you have any other questions.